OpenStudy (anonymous):

if a plane can travel 490 mph with the wind and 430mph against the wind, find the speed of the plane in still air

OpenStudy (shayaan_mustafa):

hello deano8677 :) make a diagram first.

OpenStudy (anonymous):

i made one and it doesnt seem to be helping me :(

OpenStudy (lgbasallote):

use algebra! :DDD upstream speed = speed of plane (still) - speed of current downstream speed = speed of plane (still) + speed of current need more hints? or are you good? :D

OpenStudy (anonymous):

(490+430)/2=460mph

OpenStudy (anonymous):

well ya i kinda do need more hints lol

OpenStudy (anonymous):

that was more confusing to me than the ?

OpenStudy (lgbasallote):

let x = speed of plane (still) and c = current downstream speed = 490 mph = x + c upstream = 430 = x - c x + c = 490 x - c = 430 what about now? getting it? :DDD

OpenStudy (anonymous):

so it would be 460?

OpenStudy (lgbasallote):

hmmm remember you're looking for speed of plane (still) which is x...you use systems to cancel the speed of current (which is c) so you add the system to cancel c

OpenStudy (anonymous):

so my formula is x-c(X) + c=490

OpenStudy (lgbasallote):

wait..what did you do? x + c = 490 x - c = 430 add the 2 equations (x+x) + (c - c) = (490 +430) 2x + 0 = 920 getting it now?

OpenStudy (anonymous):

im sorry my computer froze on me

OpenStudy (anonymous):

so wouldnt that make it 460?

OpenStudy (anonymous):

2x=920 x=460

OpenStudy (lgbasallote):

right! :D

OpenStudy (anonymous):

LOL i said that 20 minutes ago!!! thanks for all your help though!!!!!! greatly appreciated

OpenStudy (anonymous):

that was one of those its so easy its hard ?s

OpenStudy (lgbasallote):

ohhh...well the time thing of openstudy is buggy lol....deceived me haha wish they could fix it @_@ anyway glad to have helpedyou understand..it just takes experience haha